OpenStudy (anonymous):

The sum of the angle measures of a polygon with s sides is 2,340o. Find s

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

180*(n-2)=2340

OpenStudy (alexandervonhumboldt2):

180*(s-2)=2340

OpenStudy (alexandervonhumboldt2):

15

OpenStudy (alexandervonhumboldt2):

@bohotness

OpenStudy (anonymous):

@AlexandervonHumboldt2 did i do this right?

OpenStudy (alexandervonhumboldt2):

15 should be

OpenStudy (alexandervonhumboldt2):

sum of inteior anles of poly is 180*(n-2)

OpenStudy (cwrw238):

yes 15 is correct

OpenStudy (xapproachesinfinity):

seems good to me

OpenStudy (cwrw238):

one other bit of knowledge which might be useful sometimes is the sum of the exterior angles of ALL polygons is 360 degrees

OpenStudy (anonymous):

180-124=y x+y+125+65=360 y=54 x=114

OpenStudy (anonymous):

did i do this right?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@xapproachesinfinity @cwrw238

OpenStudy (cwrw238):

No angle y = 180 - 116 = 64

OpenStudy (cwrw238):

and angle x = 360 - (64+72+65)

OpenStudy (cwrw238):

sorry that's 125 not 65

OpenStudy (anonymous):

then solve for x?

OpenStudy (cwrw238):

yes x = 360 - (125+72 + 65)

OpenStudy (cwrw238):

the 4 angles of a quadrilateral add up to 360 degrees.

OpenStudy (cwrw238):

well you can work that out i'm sure add 125+64+72 using your calculator the subtract that from 360

OpenStudy (anonymous):

x=360 - (262) ? @cwrw238

OpenStudy (anonymous):

sooo 98

OpenStudy (cwrw238):

360 - 261

OpenStudy (cwrw238):

99

OpenStudy (cwrw238):

oh - my error - i wrote 65 instead of 64 360 - (125+72+64) = 360 - 261 = 99

OpenStudy (cwrw238):

i think im a little tired ...

OpenStudy (anonymous):

y=180 - 116 =64 x=360 - (54+72+125) x=360-261 x=99

OpenStudy (anonymous):

@cwrw238 @AlexandervonHumboldt2